OpenStudy (anonymous):

Need help integrating 1/(e^2x) -1. Can anyone help?

OpenStudy (ash2326):

Is this your question \[ \int \frac{1}{e^{2x}-1} dx\]

OpenStudy (anonymous):

Substitute u=e^2x Then it is easy

OpenStudy (ash2326):

alsunshine will you be able to solve it?

OpenStudy (anonymous):

no the dx isn't on top I'm not sure what to do

OpenStudy (ash2326):

dx will come automatically \[ \int \frac{1}{e^{2x}-1} dx\]

OpenStudy (anonymous):

i don't think so

OpenStudy (ash2326):

as we are integrating with respect to x

OpenStudy (anonymous):

i mean the du will equal 2e^2xdx. I'm not sure what to do with that.

OpenStudy (ash2326):

u= e^2x , so du=2udx can you solve it now?

OpenStudy (anonymous):

i think you have work to do. are going to have two rewrite either as \[\int\frac{1}{(e^x+1)(e^x-1)}dx\] or as \[\frac{1}{2}\int \frac{1}{e^u-1}\] and then use partial fractions

OpenStudy (anonymous):

@ash you may have something, but i don't see it. can you explain?

OpenStudy (ash2326):

We have \[ u=e^{2x} \] \[ du =2 e^{2x} dx\] but u=e^2x so \[ du = 2u dx\] Now the integral is \[ \int \frac{ 2u }{u-1} du\] this can be written as \[ \int \frac{ 2(u-1+1) }{u-1} du\] so we get \[ \int 1+\frac{ 2(1) }{u-1} du\] we get \[u + 2\ln(u-1)+c\] where u=e^2x

OpenStudy (ash2326):

sunshine you got it?

OpenStudy (anonymous):

looks good to me!

OpenStudy (anonymous):

thanks so much guys. that looks like a mess

OpenStudy (ash2326):

didn't you understand?

OpenStudy (anonymous):

understood until you rewrote as 2(u-1+1)/u-1?

OpenStudy (anonymous):

how did you get that from 2u

OpenStudy (ash2326):

Sorry I was outside we have \[2(u)\] subtract and add 1 to u \[2(u-1+1)\] This doesn't change any thing

OpenStudy (ash2326):

now \[\frac{2(u-1+1)}{u-1}=\frac{2(u-1)}{u-1}+\frac{2(1)}{u-1}=> 2+ \frac{2}{u-1}\]

OpenStudy (anonymous):

oh ok! Thanks so much for your help